This question has multiple parts. Work all the parts to get the most points. A 50.00-mL aliquot of 0.1900 M NaOH is titrated50.00 mL of acid pH = 51.00 mL of acid pH = 55.00 mL of acid pH =
Answer #1

1)when 0.0 mL of NaOH is added

Given:

M(HCl) = 0.19 M

V(HCl) = 50 mL

M(NaOH) = 0.19 M

V(NaOH) = 0 mL

mol(HCl) = M(HCl) * V(HCl)

mol(HCl) = 0.19 M * 50 mL = 9.5 mmol

mol(NaOH) = M(NaOH) * V(NaOH)

mol(NaOH) = 0.19 M * 0 mL = 0 mmol

We have:

mol(HCl) = 9.5 mmol

mol(NaOH) = 0 mmol

0 mmol of both will react

remaining mol of HCl = 9.5 mmol

Total volume = 50.0 mL

[H+]= mol of acid remaining / volume

[H+] = 9.5 mmol/50.0 mL

= 0.19 M

use:

pH = -log [H+]

= -log (0.19)

= 0.7212

Answer: 0.721

2)when 10.0 mL of NaOH is added

Given:

M(HCl) = 0.19 M

V(HCl) = 50 mL

M(NaOH) = 0.19 M

V(NaOH) = 10 mL

mol(HCl) = M(HCl) * V(HCl)

mol(HCl) = 0.19 M * 50 mL = 9.5 mmol

mol(NaOH) = M(NaOH) * V(NaOH)

mol(NaOH) = 0.19 M * 10 mL = 1.9 mmol

We have:

mol(HCl) = 9.5 mmol

mol(NaOH) = 1.9 mmol

1.9 mmol of both will react

remaining mol of HCl = 7.6 mmol

Total volume = 60.0 mL

[H+]= mol of acid remaining / volume

[H+] = 7.6 mmol/60.0 mL

= 0.1267 M

use:

pH = -log [H+]

= -log (0.1267)

= 0.8973

Answer: 0.897

3)when 40.0 mL of NaOH is added

Given:

M(HCl) = 0.19 M

V(HCl) = 50 mL

M(NaOH) = 0.19 M

V(NaOH) = 40 mL

mol(HCl) = M(HCl) * V(HCl)

mol(HCl) = 0.19 M * 50 mL = 9.5 mmol

mol(NaOH) = M(NaOH) * V(NaOH)

mol(NaOH) = 0.19 M * 40 mL = 7.6 mmol

We have:

mol(HCl) = 9.5 mmol

mol(NaOH) = 7.6 mmol

7.6 mmol of both will react

remaining mol of HCl = 1.9 mmol

Total volume = 90.0 mL

[H+]= mol of acid remaining / volume

[H+] = 1.9 mmol/90.0 mL

= 2.111*10^-2 M

use:

pH = -log [H+]

= -log (2.111*10^-2)

= 1.6755

Answer: 1.68

4)when 45.0 mL of NaOH is added

Given:

M(HCl) = 0.19 M

V(HCl) = 50 mL

M(NaOH) = 0.19 M

V(NaOH) = 45 mL

mol(HCl) = M(HCl) * V(HCl)

mol(HCl) = 0.19 M * 50 mL = 9.5 mmol

mol(NaOH) = M(NaOH) * V(NaOH)

mol(NaOH) = 0.19 M * 45 mL = 8.55 mmol

We have:

mol(HCl) = 9.5 mmol

mol(NaOH) = 8.55 mmol

8.55 mmol of both will react

remaining mol of HCl = 0.95 mmol

Total volume = 95.0 mL

[H+]= mol of acid remaining / volume

[H+] = 0.95 mmol/95.0 mL

= 1*10^-2 M

use:

pH = -log [H+]

= -log (1*10^-2)

= 2

Answer: 2.00

5)when 49.0 mL of NaOH is added

Given:

M(HCl) = 0.19 M

V(HCl) = 50 mL

M(NaOH) = 0.19 M

V(NaOH) = 49 mL

mol(HCl) = M(HCl) * V(HCl)

mol(HCl) = 0.19 M * 50 mL = 9.5 mmol

mol(NaOH) = M(NaOH) * V(NaOH)

mol(NaOH) = 0.19 M * 49 mL = 9.31 mmol

We have:

mol(HCl) = 9.5 mmol

mol(NaOH) = 9.31 mmol

9.31 mmol of both will react

remaining mol of HCl = 0.19 mmol

Total volume = 99.0 mL

[H+]= mol of acid remaining / volume

[H+] = 0.19 mmol/99.0 mL

= 1.919*10^-3 M

use:

pH = -log [H+]

= -log (1.919*10^-3)

= 2.7169

Answer: 2.72

6)when 50.0 mL of NaOH is added

Given:

M(HCl) = 0.19 M

V(HCl) = 50 mL

M(NaOH) = 0.19 M

V(NaOH) = 50 mL

mol(HCl) = M(HCl) * V(HCl)

mol(HCl) = 0.19 M * 50 mL = 9.5 mmol

mol(NaOH) = M(NaOH) * V(NaOH)

mol(NaOH) = 0.19 M * 50 mL = 9.5 mmol

We have:

mol(HCl) = 9.5 mmol

mol(NaOH) = 9.5 mmol

9.5 mmol of both will react to form neutral solution

hence pH of solution will be 7

Answer: 7.00

7)when 51.0 mL of NaOH is added

Given:

M(HCl) = 0.19 M

V(HCl) = 50 mL

M(NaOH) = 0.19 M

V(NaOH) = 51 mL

mol(HCl) = M(HCl) * V(HCl)

mol(HCl) = 0.19 M * 50 mL = 9.5 mmol

mol(NaOH) = M(NaOH) * V(NaOH)

mol(NaOH) = 0.19 M * 51 mL = 9.69 mmol

We have:

mol(HCl) = 9.5 mmol

mol(NaOH) = 9.69 mmol

9.5 mmol of both will react

remaining mol of NaOH = 0.19 mmol

Total volume = 101.0 mL

[OH-]= mol of base remaining / volume

[OH-] = 0.19 mmol/101.0 mL

= 1.881*10^-3 M

use:

pOH = -log [OH-]

= -log (1.881*10^-3)

= 2.7256

use:

PH = 14 - pOH

= 14 - 2.7256

= 11.2744

Answer: 11.27

8)when 55.0 mL of NaOH is added

Given:

M(HCl) = 0.19 M

V(HCl) = 50 mL

M(NaOH) = 0.19 M

V(NaOH) = 55 mL

mol(HCl) = M(HCl) * V(HCl)

mol(HCl) = 0.19 M * 50 mL = 9.5 mmol

mol(NaOH) = M(NaOH) * V(NaOH)

mol(NaOH) = 0.19 M * 55 mL = 10.45 mmol

We have:

mol(HCl) = 9.5 mmol

mol(NaOH) = 10.45 mmol

9.5 mmol of both will react

remaining mol of NaOH = 0.95 mmol

Total volume = 105.0 mL

[OH-]= mol of base remaining / volume

[OH-] = 0.95 mmol/105.0 mL

= 9.048*10^-3 M

use:

pOH = -log [OH-]

= -log (9.048*10^-3)

= 2.0435

use:

PH = 14 - pOH

= 14 - 2.0435

= 11.9565

Answer: 11.96